OpenStudy (anonymous):

Sally found that a solution of x = 2 was extraneous for her function. What would be a possible denominator in Sally's function? 4x + 8 2x + 12 x2 - 4 x2 - 16

OpenStudy (anonymous):

@kritikachandak

zepdrix (zepdrix):

extraneous is bad. example: if x=3 is an extraneous solution, we might have a denominator of:\(\Large\rm x-3\)

OpenStudy (anonymous):

extraneous means its not relevant to the equation right

zepdrix (zepdrix):

This denominator gives us a zero in the denominator when x=3. So it's bad!:O yah, it's not a workable solution. we can ignore it as a solution.

zepdrix (zepdrix):

Since we have x=2 as extraneous, (x-2) is the form we're looking for in the denominator. Do any of our options give us that? We'll have to factor to see correctly.

OpenStudy (anonymous):

ok so how do we find the denominator?

zepdrix (zepdrix):

x=2 is extraneous, \(\Large\rm \color{red}{(x-2)}\) is the factor we'e looking for. \(\Large\rm 4x+8=4(x+2)\) Hmm that one didn't give us a factor of x-2. So it's not option A.

zepdrix (zepdrix):

Understand how that works? :o

OpenStudy (anonymous):

yessir

zepdrix (zepdrix):

ahh sorry class is starting :c i gotta go

OpenStudy (anonymous):

:( darnnit

zepdrix (zepdrix):

\(\Large\rm x^2-4=(x+2)\color{red}{(x-2)}\)

OpenStudy (anonymous):

is it x^2-4 real quick?

zepdrix (zepdrix):

Remember how to factor the difference of squares?

zepdrix (zepdrix):

oh boy the website is tweaking out on me :c